please help me with this..
The length of a rectangular field is 2 yards more than its width. Find the width if the area of the field is 120 yd^2
thank you so much!!

Answer by tiger(2)

let width=x
so, length=x+2
so, area will be x(x+2)=120
x^2 +2x-120=0
x^2 +12x -10x -120 -0
x(x+12)-10(x+12)=0
so x=-12 or 10
so answer will be 10.....